You have a solution that is 50% alcohol and a solution that is 80% alcohol. How much of each do you need to mix to get 3 gallons of a solution that is 60% alcohol?
Answer by nerdybill(7384)

You have a solution that is 50% alcohol and a solution that is 80% alcohol. How much of each do you need to mix to get 3 gallons of a solution that is 60% alcohol?
.
Let x = amount of 50% alcohol
then
3-x = amount of 80% alcohol
.
.50x + .80(3-x) = .60(3)
.50x + 2.4 - .80x = 1.8
2.4 -.30x = 1.8
-.30x = -0.6
x = 2 gallons (50% alcohol)
.
80% alcohol:
3-x = 3-2 = 1 gallon (80% alcohol)